# Sample Compression Unleashed: New Generalization Bounds for Real Valued Losses

### v2

@@ 4.4 Amazon polarity
Finally, we train DistilBERT (Sanh et al.,, 2020) on the Amazon reviews polarity dataset (Zhang et al.,, 2015). Using P2L, we fine-tune the pretrained language model on 10% of the dataset, for a total of 360k datapoints, and evaluate the model on the test set, which comprises 400k datapoints. As the model and dataset are quite large, we run the experiments for each hyperparameter combination only once. We pre-train the model on half of the training dataset and then use P2L on the other half of the training set. We add 32 datapoints at a time in the compression set and early stop the training of the model if its validation loss has not decreased for 20 epochs. In this experiment, we study our new $\mathrm{kl}$ bound on the zero-one loss and on the bounded cross-entropy loss. Moreover, we compute the binomial approximation bound of Corollary 9. The P2L bound (Theorem 2) is invalid in this setting, as the model doesn’t reach zero errors. The PAC-Bayesian bound of Theorem 10 could be computed on both metrics, but it would necessitate to train 132M parameters (twice the number of parameters of DistilBERT). Many new generalization bounds and approaches were presented for very large models (Lotfi et al., 2024a, ; Lotfi et al., 2024b, ; Zekri et al.,, 2024; Su et al.,, 2024), such as large language models. However, most approaches are not suited for classification and regression, as they are derived for language modeling objectives.
We present the results in Table 5. First of all, we observe that training the model using P2L only incurs a loss of about a percent for the train, validation and test error. It achieves this error whilst being trained on about 1% of the dataset, as the compression set size is 1472 and the training set size is 144k. Both for the error and the cross-entropy loss, the bound is tight and non-vacuous. Our bound is much tighter than the binomial approximation bound, with a certificate of $14.67\%$ for a train error of $4.25\%$. Despite the 66M parameters of DistilBERT, we are able to obtain tight generalization guarantees by simply changing the training loop of the model.
@@ 5 Conclusion
We developed novel generalization bounds for real-valued losses and sample-compressed predictors. These bounds leverage the comparator functions studied in the PAC-Bayes theory. We provide results for bounded and unbounded losses, under different assumptions. We empirically verified the tightness of the proposed bounds, showing that it is almost as tight as the binomial tail inversion, which, however, holds only for a less general setting. We trained neural networks with 66M parameters and obtained tight guarantees, without suffering from the cost of the number of parameters. This highlights an important asset of the sample compression framework: Two models achieving the same empirical loss using the same amount of datapoints (compression set size) share the same guarantees (bound value), regardless of their size in terms of the number of trainable parameters.
In future works, we could leverage the possibility of having a message in the compression scheme, by training models such as the set covering machine (Laviolette et al.,, 2005) or decision trees (Shah,, 2007), which both use binary sequences to specify how to reconstruct the model. Finally, although P2L is generally able to train good performing models, it is unclear that its sample selection heuristic is optimal for neural networks. Trying different heuristics, e.g., that optimize for sample diversity, could lead to improved performance and guarantees for the models.
Acknowledgments
@@ Appendix A Experiments
The experiments were run on two different devices. The experiments with PBB algorithm and the regression datasets were run on Python 3.12.2 on a computer with a NVIDIA GeForce RTX 4090. The experiments on MNIST were run on Python 3.12.3 on a computer with a NVIDIA GeForce RTX 2080 Ti. The libraries used for each environment can be found with the code. Notably, we use PyTorch (Ansel et al.,, 2024) (BSD 3-Clause License), Lightning (Falcon and The PyTorch Lightning team,, 2019) (Apache 2.0 license), Weight and Biases (Biewald,, 2020) (MIT License), Scikit-Learn (Pedregosa et al.,, 2011) (BSD 3-Clause License), NumPy (Harris et al.,, 2020) (NumPy license) and Transformer (Wolf et al.,, 2020) (Apache 2.0 license). For all experiments, we run the code with the following seeds : $[1,2,3,4,42]$.
We give information on the datasets used in the experiments.
For the classification problems, we use the MNIST dataset (LeCun et al.,, 1998) (MIT License) and the amazon polarity dataset (Zhang et al.,, 2015) (Apache 2.0 License). All MNIST derived-dataset are composed of 784 real-valued features. For the multi-class classification problems on MNIST, we denote MNIST ($p$%) to say that we pre-train the model on $p$of the data, where $p$ is a hyperparameter. For the Amazon polarity dataset, we chose 10% of the dataset to create a 360000 datapoints dataset. We then use 50% to pre-train the model and split the rest into a training and validation set. The datapoints are textual reviews and the labels are binary. The descriptions of the dataset are presented in Table 6.
@@
Amazon Polarity 180000 144000 36000 400000
For the regression problems, we train our models on five dataset : the Combined Cycle Power Plant (Tüfekci,, 2014; Tfekci and Kaya,, 2014), the Infrared Thermography Temperature (Wang et al.,, 2021, 2023), the Airfoil Self-Noise (Brooks et al., 1989b, ; Brooks et al., 1989a, ), the Parkinsons Telemonitoring (Tsanas et al.,, 2009; Tsanas and Little,, 2009), the Concrete Compressive Strength (Yeh, 1998b, ; Yeh, 1998a, ). The descriptions of the dataset are presented in Table 7. All datasets were chosen from the UCI dataset repository. Powerplant, Airfoil, Parkinson and Concrete are under the CC-BY 4.0 license. The Infrared dataset is under the CC0 license.
@@ A.1.1 Binary MNIST problems
For the binary MNIST problems, we used the following hyperparameters.
  • Model type : [MLP, CNN]
  • Dropout probability : [0.1, 0.2]
  • Training learning rate : $[1\textrm{e}-2,1\textrm{e}-3,5\textrm{e}-3,1\textrm{e}-4]$
@@
At each iteration, the P2L algorithm adds one datapoint to the compression set.
@@
In the following tables, we present the results for the MLP, both trained fully using P2L and early-stopped, respectively in Table 8 and in Table 9. Moreover, in Fig. 2, we present the results not present in Fig. 1.
Table 8: Results for the MLPs trained using P2L on the binary MNIST problems. The results displayed obtained the tightest P2L bound. All metrics presented are in percents (%), with the exception of $\operatorname{|\mathbf{i}|}$.
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound P2L bound $\operatorname{|\mathbf{i}|}$ Baseline test error
MNIST08 0.41$\pm$0.14 0.40$\pm$0.08 6.56$\pm$0.30 6.51$\pm$0.30 1.42$\pm$0.08 128.2$\pm$7.4 0.34$\pm$0.07
MNIST17 0.37$\pm$0.14 0.47$\pm$0.17 4.93$\pm$0.27 4.89$\pm$0.27 1.01$\pm$0.07 99.0$\pm$7.0 0.33$\pm$0.08
MNIST23 0.87$\pm$0.24 0.58$\pm$0.12 12.21$\pm$0.29 12.17$\pm$0.29 3.06$\pm$0.09 296.6$\pm$9.4 0.36$\pm$0.14
MNIST49 1.19$\pm$0.33 1.04$\pm$0.10 14.41$\pm$0.05 14.37$\pm$0.05 3.78$\pm$0.02 361.4$\pm$1.9 0.96$\pm$0.14
MNIST56 0.68$\pm$0.17 0.65$\pm$0.05 10.35$\pm$0.31 10.30$\pm$0.31 2.48$\pm$0.09 223.0$\pm$8.9 0.59$\pm$0.15
Table 9: Results for the MLPs trained using P2L on the binary MNIST problems and stopped at the iteration with the minimum $\mathrm{kl}$ bound. The results displayed obtained the tightest $\mathrm{kl}$ bound. All metrics presented are in percents (%), with the exception of $\operatorname{|\mathbf{i}|}$.
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound Train error $\operatorname{|\mathbf{i}|}$ Baseline test error
MNIST08 1.11$\pm$0.52 1.04$\pm$0.67 5.46$\pm$0.53 7.77$\pm$1.64 0.85$\pm$0.71 59.2$\pm$34.4 0.34$\pm$0.07
MNIST17 0.88$\pm$0.39 0.80$\pm$0.29 4.02$\pm$0.36 5.49$\pm$0.77 0.50$\pm$0.26 44.0$\pm$15.0 0.33$\pm$0.08
MNIST23 1.93$\pm$0.49 1.59$\pm$0.43 10.86$\pm$0.19 13.23$\pm$0.74 1.27$\pm$0.41 146.0$\pm$25.8 0.36$\pm$0.14
MNIST49 2.28$\pm$0.53 2.07$\pm$0.58 13.14$\pm$0.32 15.08$\pm$0.99 1.22$\pm$0.47 202.0$\pm$30.6 0.96$\pm$0.14
MNIST56 1.97$\pm$0.53 1.88$\pm$0.44 8.85$\pm$0.58 11.78$\pm$1.44 1.38$\pm$0.61 92.0$\pm$27.9 0.59$\pm$0.15
@@
Figure 2: Illustration of the behavior of the $\mathrm{kl}$ bound throughout P2L iterations for the five different seeds of the hyperparameter combination that achieved the minimal P2L bound. We mark the minimal $\mathrm{kl}$ bound for each seed with a diamond ($\blacklozenge$).
@@ A.1.2 MNIST problems
Table 10: Train metrics on MNIST (risk 01) (in percents)
Model Train error Test Error $\mathrm{kl}$ bound Binomial bound Compression set size Baseline test error
P2L 0.0$\pm$0.0 1.06$\pm$0.10 6.15$\pm$ 1.51 6.13$\pm$1.51 275.20$\pm$ 82.46 0.0108$\pm$0.0010
@@
  • Size of pretraining set : $[10\%,20\%,50\%]$
  • Pretraining epochs : [50, 100]
  • Pretraining learning rate : $[1\textrm{e}-2,1\textrm{e}-3,1\textrm{e}-4]$
  • Dropout probability : [0.1, 0.2]
  • Training learning rate : $[1\textrm{e}-3,5\textrm{e}-3,1\textrm{e}-4]$
At each iteration, the P2L algorithm adds 32 datapoint to the compression set. To compute bounds for the cross-entropy loss, we clamp the log-probabilities to be greater or equal than $\ln(10^{-5})$ (Pérez-Ortiz et al.,, 2021; Dziugaite and Roy,, 2018), as follows :
@@
where $h(\boldsymbol{x})=\quantity(h(\boldsymbol{x})_{1},\ldots,h(\boldsymbol{x})_{C})$ is the output of the neural network and $C$ is the number of classes. The loss then takes values between $[0,-\ln(10^{-5})]$. We will use the same bounded cross-entropy loss for the following experiments.
@@
For the PAC-Bayes with Backprop (PBB) algorithm, we used the GitHub repository associated to the article of Pérez-Ortiz et al., (2021). We used the hyperparameter grid proposed by the article, with the exception of the dropout, which we kept similar to the other experiments. We used $\delta=\delta^{\prime}=0.01$ to compute Theorem 10. We used $m=5000$ Monte Carlo sampling instead of the value $m=150000$ found in the code, as it takes several hours to run.
  • Scale parameter of the prior distribution : $[0.1,0.05,0.04,0.03,0.02,0.01,0.005]$
  • Training learning rate : $[1e-3,5e-3,1e-2]$
  • Pre-training learning rate : $[1e-3,5e-3,1e-2]$
  • Momentum : $[0.95,0.99]$
  • Dropout probability : $[0.1,0.2]$

### v3

@@ 4.4 Amazon polarity
Finally, we train DistilBERT (Sanh et al.,, 2019) on the Amazon reviews polarity dataset (Zhang et al.,, 2015). Using P2L, we fine-tune the pretrained language model on 10% of the dataset, for a total of 360k datapoints, and evaluate the model on the test set, which comprises 400k datapoints. We pre-train the model on half of the training dataset and then use P2L on the other half of the training set. We add 32 datapoints at a time in the compression set and early stop the training of the model if its validation loss has not decreased for 20 epochs. In this experiment, we study our new $\mathrm{kl}$ bound on the zero-one loss and on the bounded cross-entropy loss. Moreover, we compute the binomial approximation bound of Corollary 10. The P2L bound (Theorem 2) is inapplicable in this setting, as the model doesn’t reach zero errors. The PAC-Bayesian bound of Theorem 11 could be computed on both metrics, but it would necessitate to train 132M parameters (twice the number of parameters of DistilBERT). Many new generalization bounds and approaches were presented for very large models (Lotfi et al.,, 2024, 2025; Zekri et al.,, 2024; Su et al.,, 2025), such as large language models. However, most approaches are not suited for classification and regression, as they are derived for language modeling objectives.
From the results displayed by Table 6, we first observe that training the model using P2L only incurs a loss of about a percent for the train, validation and test error compared to the baseline. It achieves this error whilst being trained on about 0.8% of the dataset, as the compression set size is roughly 1138 datapoints and the training set size is 144k. Both for the error and the cross-entropy loss, the bound is tight and non-vacuous. Our bound is much tighter than the binomial approximation bound, with a certificate of $13.91\%$ for a train error of $4.73\%$. Thus, despite the 66M parameters of DistilBERT, we are able to obtain tight generalization guarantees by simply changing the training loop of the model for the P2L scheme.
@@ 5 CONCLUSION
We proposed novel generalization bounds for real-valued losses and sample-compressed predictors. These bounds leverage the comparator functions studied in the PAC-Bayes theory. We provide results for bounded and unbounded losses, under different assumptions. We empirically verified the tightness of the proposed bounds, showing that it is almost as tight as the binomial tail inversion, which, however, holds only for a less general setting. We trained neural networks with 66M parameters and obtained tight guarantees, without suffering from the cost of the number of parameters. This highlights an important asset of the sample compression framework: Two models achieving the same empirical loss using the same amount of datapoints (compression set size) share the same guarantees (bound value), regardless of their size in terms of the number of trainable parameters.
In future works, we could leverage the possibility of having a message in the compression scheme, by training models such as the set covering machine (Laviolette et al.,, 2005) or decision trees (Shah,, 2007), which both use binary sequences to specify how to reconstruct the model. Finally, although P2L is generally able to train good performing models, it is unclear that its sample selection heuristic is optimal for neural networks. Trying different heuristics, e.g., that optimize for sample diversity, could lead to further improvements.
Acknowledgements
We would like to thank the anonymous reviewers for their helpful comments. We also wish to thank Benjamin Leblanc and Sokhna Diarra Mbacke for the insightful discussions and their help proof-reading the manuscript.
@@ Acknowledgements
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes, in Section 2.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Not applicable.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes, the code can be found here : https://github.com/GRAAL-Research/pick-to-learn.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes, the only assumption is that the data is i.i.d. and is mentioned in our setting in Section 2.
    2. (b)
      Complete proofs of all theoretical results. Yes, in Appendix C.
    3. (c)
      Clear explanations of any assumptions. The only assumption is that the data is i.i.d., which is the classical setting in the literature.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). All the code and instruction can be found in the repository. The data is imported directly when the code is executed.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). All the training details can be found in Appendix A and the code.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). All metrics are defined in either Section 4 or Appendix A. We present the mean and standard deviation over five seeds.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes, in Appendix A.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes.
    2. (b)
      The license information of the assets, if applicable. Yes, in Appendix A.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Yes, the code can be found here : https://github.com/GRAAL-Research/pick-to-learn.
    4. (d)
      Information about consent from data providers/curators. Not applicable.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not applicable.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not applicable
Sample compression unleashed :
@@ Appendix A EXPERIMENTS
Devices.
The experiments were run on two different devices. The experiments with PBB algorithm and the regression datasets were run on Python 3.12.2 on a computer with a NVIDIA GeForce RTX 4090. The experiments on MNIST were run on Python 3.12.3 on a computer with a NVIDIA GeForce RTX 2080 Ti.
Librairies.
The libraries used for each environment can be found with the code. Notably, we use PyTorch (Ansel et al.,, 2024) (BSD 3-Clause License), Lightning (Falcon and The PyTorch Lightning team,, 2019) (Apache 2.0 license), Weight and Biases (Biewald,, 2020) (MIT License), Scikit-Learn (Pedregosa et al.,, 2011) (BSD 3-Clause License), NumPy (Harris et al.,, 2020) (NumPy license) and Transformer (Wolf et al.,, 2020) (Apache 2.0 license). For all experiments, we run the code with the following seeds : $\{1,2,3,4,42\}$.
Datasets.
For the classification problems, we use the MNIST dataset (LeCun et al.,, 1998) (MIT License) and the amazon polarity dataset (Zhang et al.,, 2015) (Apache 2.0 License). All MNIST derived-dataset are composed of 784 real-valued features. For the multi-class classification problems on MNIST, we denote MNIST ($p$%) to say that we pre-train the model on $p$of the data, where $p$ is a hyperparameter. For the Amazon polarity dataset, we chose 10% of the dataset to create a 360k datapoints dataset. We then use 50% to pre-train the model and split the rest into a training and validation set. The datapoints are textual reviews and the labels are binary. The description of the dataset are presented in Table 7.
@@
For the regression problems, we train our models on five datasets : the Combined Cycle Power Plant (Tüfekci,, 2014; Tfekci and Kaya,, 2014), the Infrared Thermography Temperature (Wang et al.,, 2021, 2023), the Airfoil Self-Noise (Brooks et al., 1989b, ; Brooks et al., 1989a, ), the Parkinsons Telemonitoring (Tsanas et al.,, 2009; Tsanas and Little,, 2009) and the Concrete Compressive Strength (Yeh, 1998b, ; Yeh, 1998a, ). The descriptions of the dataset are presented in Table 8. All datasets were chosen from the UCI dataset repository. Powerplant, Airfoil, Parkinson and Concrete are under the CC-BY 4.0 license. The Infrared dataset is under the CC0 license.
@@ A.1.1 Binary MNIST problems
For the binary MNIST problems, we used the following hyperparameters for both MLP and CNN architectures.
  • Dropout probability : $\{0.1,0.2\}$
  • Training learning rate : $\{10^{-2},10^{-3},5\times 10^{-3},10^{-4}\}$
@@
At each iteration, the P2L algorithm adds one datapoint to the compression set. We use $h_{0}$ a randomly initialized neural network.
@@
We present the results for the MLP, both trained fully using P2L and early-stopped, respectively in Table 9 and in Table 10. Moreover, Fig. 2 displays the results not present in Fig. 1.
Table 9: Results for the MLPs trained using P2L on the binary MNIST problems. The results displayed obtained the tightest P2L bound. All metrics presented are in percents (%).
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound P2L bound $\operatorname{|\mathbf{i}|}/n$ Baseline test error
MNIST08 0.41$\pm$0.14 0.40$\pm$0.08 6.56$\pm$0.30 6.51$\pm$0.30 1.42$\pm$0.08 1.21$\pm$0.07 0.34$\pm$0.07
MNIST17 0.37$\pm$0.14 0.47$\pm$0.17 4.93$\pm$0.27 4.89$\pm$0.27 1.01$\pm$0.07 0.85$\pm$0.06 0.33$\pm$0.09
MNIST23 0.87$\pm$0.24 0.58$\pm$0.12 12.21$\pm$0.29 12.17$\pm$0.29 3.06$\pm$0.09 2.73$\pm$0.09 0.36$\pm$0.14
MNIST49 1.19$\pm$0.33 1.04$\pm$0.10 14.41$\pm$0.05 14.37$\pm$0.05 3.78$\pm$0.02 3.41$\pm$0.02 0.96$\pm$0.15
MNIST56 0.68$\pm$0.17 0.65$\pm$0.05 10.35$\pm$0.31 10.30$\pm$0.31 2.48$\pm$0.09 2.18$\pm$0.09 0.59$\pm$0.01
Table 10: Results for the MLPs trained using P2L on the binary MNIST problems and stopped at the iteration with the minimum $\mathrm{kl}$ bound. The results displayed obtained the tightest $\mathrm{kl}$ bound. All metrics presented are in percents (%).
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound Train error $\operatorname{|\mathbf{i}|}/n$ Baseline test error
MNIST08 1.11$\pm$0.52 1.04$\pm$0.67 5.46$\pm$0.53 7.77$\pm$1.64 0.85$\pm$0.71 0.56$\pm$0.32 0.34$\pm$0.07
MNIST17 0.88$\pm$0.39 0.80$\pm$0.29 4.02$\pm$0.36 5.49$\pm$0.77 0.50$\pm$0.26 0.38$\pm$0.13 0.33$\pm$0.09
MNIST23 1.93$\pm$0.49 1.59$\pm$0.43 10.86$\pm$0.19 13.23$\pm$0.74 1.27$\pm$0.41 1.34$\pm$0.24 0.36$\pm$0.14
MNIST49 2.28$\pm$0.53 2.07$\pm$0.58 13.14$\pm$0.32 15.08$\pm$0.99 1.22$\pm$0.47 1.90$\pm$0.29 0.96$\pm$0.15
MNIST56 1.97$\pm$0.53 1.88$\pm$0.44 8.85$\pm$0.58 11.78$\pm$1.44 1.38$\pm$0.61 0.90$\pm$0.27 0.59$\pm$0.01
@@
Figure 2: Illustration of the behavior of the $\mathrm{kl}$ bound throughout P2L iterations for the five different random seed initializations sharing the hyperparameter combination that achieved the minimal averaged P2L bound. The diamonds ($\blacklozenge$) mark the minimal $\mathrm{kl}$ bound for each run.
@@ A.1.2 MNIST problems
Table 11: Classification risk achieved by the CNNs on MNIST. The results displayed obtained the smallest $\mathrm{kl}$ bound. All metrics are presented in percents (%).
Learning algorithm Train error Test Error $\mathrm{kl}$ bound Binomial bound $\operatorname{|\mathbf{i}|}/n$ Baseline test error
P2L 0.0$\pm$0.0 1.14$\pm$0.07 7.11$\pm$0.22 7.08$\pm$0.22 1.15$\pm$0.34 1.08$\pm$0.09
@@
  • Size of pretraining set : $\{10\%,20\%,50\%\}$
  • Pretraining epochs : $\{50,100\}$
  • Pretraining learning rate : $\{10^{-2},10^{-3},10^{-4}\}$
  • Dropout probability : $\{0.1,0.2\}$
  • Training learning rate : $\{10^{-2},5\times 10^{-3},10^{-4}\}$
At each iteration, the P2L algorithm adds 32 datapoints to the compression set. We use $h_{0}$ a neural network pre-trained on $p\%$ of the dataset (see size of pretraining set for the different values of $p$).
To compute bounds for the cross-entropy loss, we clamp the log-probabilities to be greater or equal than $\ln(10^{-5})$ (Pérez-Ortiz et al.,, 2021; Dziugaite and Roy,, 2018), as follows :
@@
where $h(\boldsymbol{x})=\quantity(h(\boldsymbol{x})_{1},\ldots,h(\boldsymbol{x})_{C})$ is the output of the neural network and $C$ is the number of classes. The loss then takes values in $[0,-\ln(10^{-5})]$. We use the same bounded cross-entropy loss for the following experiments.
@@
For the PAC-Bayes with Backprop (PBB) algorithm, we used the code of the GitHub repository provided alongside Pérez-Ortiz et al., (2021), with the same hyperparameter grid proposed by the authors, [^4] https://github.com/mperezortiz/PBB except for the dropout rate, which we kept the same as the other experiments:
  • Scale parameter of the prior distribution : $\{0.1,0.05,0.04,0.03,0.02,0.01,0.005\}$
  • Training learning rate : $\{10^{-3},5\times 10^{-3},10^{-2}\}$
  • Pre-training learning rate : $\{10^{-3},5\times 10^{-3},10^{-2}\}$
  • Momentum : $\{0.95,0.99\}$
  • Dropout probability : $\{0.1,0.2\}$
We fixed $\delta=\delta^{\prime}=0.01$ to compute Theorem 11, and performed $m=5000$ Monte Carlo sampling steps instead of using the value $m=150000$ found in the code, as it takes several hours to run.